### v2

Abstract
As agent systems powered by large language models (LLMs) advance, improving the task performance of an autonomous agent, especially in context understanding, tool usage, and response generation, has become increasingly critical. Although prior studies have advanced the overall design of LLM-based agents, systematic optimization of their internal reasoning and tool-use pipelines remains underexplored. This paper introduces an agent framework grounded in real-world practical experience, with three key innovations: (1) an adaptive prompt generation strategy that aligns with the agent’s state and task goals to improve reliability and robustness; (2) a context-aware tool orchestration module that performs tool categorization, semantic retrieval, and adaptive invocation based on user intent and context; and (3) a layered memory mechanism that integrates session memory, task history, and external summaries to improve relevance and efficiency through dynamic summarization and compression. An end-to-end framework named Jenius-Agent has been integrated with three key optimizations, including tools based on the Model Context Protocol (MCP), file input/output (I/O), and execution feedback. The experiments show a 20 percent improvement in task accuracy, along with a reduced token cost, response latency, and invocation failures. The framework is already deployed in Jenius (https://www.jenius.cn), providing a lightweight and scalable solution for robust, protocol-compatible autonomous agents.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
In the context of the increasing capability of large language models (LLMs), LLM-based autonomous agents have become a new paradigm for AI applications Yao et al. (2023); Xi et al. (2025); Wang et al. (2024). These agents are capable of understanding instructions, invoking tools, reasoning and planning, and performing complex tasks, and are widely used in a variety of fields such as research assistants, process automation, retrieval-augmented generation, and code generation and debugging Dubey et al. (2024); Masterman et al. (2024); Qin et al. (2024); Singh et al. (2025). Although current intelligent agent systems (e.g., AutoGPT, LangChain Agents, BabyAGI)  Yang et al. (2023); Topsakal and Akinci (2023); Nakajima (2023) have begun to take shape, there are still numerous challenges in their generality, stability, and manageability, especially in task precision, response reliability, and system stability in situations with many tasks Cemri et al. (2025).
Most existing agent systems rely on fixed prompts and predefined tool-use workflows, limiting their ability to understand task intent, select tools dynamically, and manage context effectively e Aquino et al. (2025); Qin et al. (2024); Schick et al. (2023). Numerous works have shown that dynamic prompt, tool retrieval, and memory management play a key role in optimizing agent execution. For instance, MCP-Zero Fei et al. (2025) enables active tool discovery for unseen tasks, while multi-agent designs Zhou et al. (2025) enhance planning and coordination through optimized prompts and collaboration strategies. The Model Context Protocol (MCP) Hou et al. (2025) standardizes context exchange, ensuring reliable tool invocation across systems. Together, these advances underscore that robust prompting, intelligent tool access, and effective memory are essential for high-quality agent execution.
Despite recent advances, autonomous agent pipelines still face three key challenges. First, fixed or generic prompts often misinterpret user intent and fail to adapt to evolving task states, causing unstable behavior and inconsistent outputs. Second, static tool lists or handcrafted rules cannot reliably choose the right tools under ambiguity or across domains, leading to unnecessary or incorrect calls. Finally, long dialogues accumulate redundant context, increasing token cost and diluting salient signals, which weakens reasoning quality.
This study starts from the basic execution process of autonomous agents and introduces three complementary optimization modules addressing the above issues:
  • Task Understanding and Prompt Optimization: Structured intent recognition was combined with refined system prompts and templates to adapt instructions to current state and goals, reducing misinterpretation and stabilizing task alignment.
  • Tool Retrieval: Dynamic retrieval and adaptive tool access were used to match user intent with context-relevant tools and to handle ambiguous user requests.
  • Hierarchical Memory Management: Redundant dialogue history was pruned to control token length, preserving essential semantics and stabilizing reasoning in long-horizon interactions.
Rather than evaluating isolated modules, this paper builds a unified framework named Jenius-Agent, which improves task accuracy, efficiency, and contextual robustness, aligned with emerging agent communication protocols (e.g., MCP, ACP, A2A) IBM BeeAI (2024); Google (2024).
The main contributions are summarized as follows:
  • We proposed a modular optimization framework integrating adaptive prompting, context-aware tool orchestration, and hierarchical memory management to mitigate context noise, tool misuse, and harmful prompting;
  • We improved task grounding and execution robustness through adaptive prompt generation and context-aware tool usage, while introducing preliminary defense strategies to strengthen the system’s resilience to malicious or erroneous inputs;
  • We designed a comprehensive evaluation framework covering the procedural, semantic, and efficiency dimensions;
  • We also conducted extensive experiments both public and real-world datasets, demonstrating consistent gains in task accuracy, response quality, token efficiency.
Figure 1: A typical ReAct style autonomous agent workflow.
2 Related Work
With the rapid advancement of LLMs, research on agent systems has expanded in multiple directions. A typical ReAct Yao et al. (2022) style autonomous agent workflow, illustrated in Figure 1, follows an interleaved cycle of reasoning, action, and feedback. In this paradigm, the LLM is no longer a passive text generator but serves as a central planner. This iterative planner–tool–environment interaction highlights three optimization focuses:(i) prompt optimization for reliable reasoning, (ii) tool selection for precise external interactions, and (iii) memory management to preserve context and limit token usage.
2.1 Prompt Engineering and Optimization
DSPy Khattab et al. provides a modular framework for prompt compilation, improving clarity and debuggability, while Prompt Engineering in LLMs Marvin et al. (2023) summarizes practical design principles. Reflect-Retry-Reward Bensal et al. (2025) further introduces adaptive optimization through self-reflection and reinforcement learning. However, current approaches remain unreliable in tool-based or long-horizon tasks, as they often depend on static templates, lack systematic evaluation, and struggle to control structured output or tool invocation consistency.
2.2 Tool Selection and Invocation
Recent efforts such as MCP-Zero Fei et al. (2025) explore adaptive tool discovery using semantic embeddings, and BioMedTools Liu et al. (2025) illustrates domain-specific coordination via curated tool metadata.However, prior approaches frequently assume clean tool descriptions and fixed invocation formats, whereas real-world systems face noisy metadata, semantic mismatches, and poorly handled runtime errors, resulting in unstable tool retrieval and parameter usage.
Figure 2: Jenius Agent Framework. The LLM acts as the central orchestrator, coordinating task execution with three core modules: adaptive prompt generation, tool retrieval, and memory management that enhance adaptability and efficiency.
2.3 Memory Management
StateFlow Wu et al. and ReAct Yao et al. (2023) demonstrate state-aware and reasoning-aligned memory design, while Recursively Summarizing Wang et al. (2025) introduces hierarchical compression for long-term coherence.Existing methods often use simple window-based truncation or stage-level summaries, risking loss of critical early context or fine-grained tool dependencies, lacking tight integration between reasoning and memory retention.
3 Method
To address the above limitations, we proposed a feedback-driven Jenius-agent(Figure 2), which integrates three coordinated optimization modules. First, adaptive prompt generation creates system prompts by combining role instructions, task state, and user context. Second, tool retrieval selects relevant tools from heterogeneous sources. Third, memory management summarizes historical interactions into compact, semantically enriched representations.
The LLM produces reasoning traces and function calls, executed by the tool execution module, with results are fed back into the agent. This loop architecture enables continuous integration of prompt adaptation, tool utilization, and memory compression, enhancing both robustness and task alignment.
The following subsections detail the design, motivation, and role of each module.
3.1 Adaptive Prompt Generation
Figure 3 illustrates the prompt generation pipeline, which derives adaptive prompts via the fusion of core behavioral constraints on agent, task-driven instructional augmentation, and history-aware adjustments.
(1) Core behavioral constraints: Defines the agent’s role, interaction protocol, response style, and operational limits to ensure consistent, safe, and task-appropriate behavior.
Figure 3: Adaptive Prompt Generation.
  • Role specification and linguistic alignment: Defines the agent’s identity and ensures linguistic coherence by aligning the response language with the user’s input, thereby avoiding unintended multilingual mixing in generated responses;
  • Intent taxonomy and routing: Classifies user queries into four semantic categories: (i) social interaction (e.g., greetings, small talk), (ii) creative generation (e.g., storytelling, content drafting), (iii) factual recall (queries resolvable via internal knowledge), and (iv) tool-augmented reasoning, further subdivided into single-tool invocation or multi-step, multi-tool orchestration. This taxonomy enables the dynamic prompt generator to tailor role instructions, reasoning strategies, and output formats according to the detected intent category, thereby ensuring that the generated responses are more precise and contextually appropriate;
  • Controlled tool invocation policy: Implements strict controls over external tool usage to ensure factual accuracy and operational reliability. The agent validates tool eligibility and parameter consistency before invocation, rejecting speculative or unverifiable requests. This mechanism mitigates hallucinations of fabricated entities (e.g., dates, names, locations), prevents redundant or cyclic calls, and enforces deterministic, auditable execution aligned with reliability standards;
  • Response structuring protocol: Determines the output format based on task semantics and usability requirements, such as structured (e.g., tables, JSON, code) or unstructured (free text).
  • Safety and content moderation: Establishes multi-layered guardrails combining hard rule-based filters and soft model-level moderation. These constraints detect and block unsafe, unethical, or privacy-violating outputs while discouraging implicit bias and emotionally manipulative expressions. The moderation layer further curbs hallucinated or contextually inappropriate content that may arise from ambiguous prompts or overgeneralized reasoning.
(2) Runtime prompt adaptation: This component dynamically adapts prompt based on task semantics, tool availability, and interaction context to ensure output alignment with user intent. It follows three principles:
  • Context-aware adaptation: Prompts are conditioned on the current query, task category, and operational constraints (e.g., tool eligibility, safety rules), guiding appropriate reasoning strategies and response formats.
  • Tool orchestration guidance: Conditioning signals direct correct and minimal invocation behavior, suppressing redundant, invalid, or hallucinated calls to improve execution reliability.
  • Fine-grained intent discrimination: The agent distinguishes between semantically similar but functionally distinct tasks (e.g., textual vs slide report) to ensure fidelity to user intent.
By integrating these mechanisms, this component complements the static behavioral constraints, improving task accuracy, reducing erroneous tool usage, and maintaining consistency and relevance across conversational turns.
(3) History-aware adaptation: Rather than relying on raw conversation logs, we employs a relevance-driven compression strategy to filter logs, only preserve key cross-turn dependencies. The resulting history is summarized and appended to the prompt, facilitating precise and coherent action planning. For history filtering details, please refer to section of ”Hierarchical Memory Management”.
3.2 Context-Aware Tool Orchestration
This module is designed to enhance the efficiency of tool access by dynamically selecting appropriate external tools based on the identified task intent. Our framework establishes a structured MCP tool management mechanism, enabling systematic optimization, extensibility, and constraint enforcement over available tools. For instance, third-party MCP tools with inadequate functional descriptions are repackaged and augmented to improve their interpretability and usability. Furthermore, tools are classified according to their functional categories, such as file management, information retrieval, image generation, and data analysis, thereby supporting organized indexing and facilitating efficient, context-aware tool discovery.
All tools are represented as high-dimensional embeddings using the Qwen3 Embedding model Zhang et al. (2025), and tool selection is formalized as a semantic similarity ranking process. The similarity between the query and each tool embedding determines their conceptual proximity, which serves as the basis for ranking and subsequent relevance filtering.
Step 1: Top-$\boldsymbol{M}$ candidate retrieval.
Candidate tools are ranked by similarity to the query embedding. The top-$M$ tools with the highest similarity values are selected as the initial candidate set, forming a shortlist for further refinement.
Step 2: Inflection point-based filtering.
To identify a meaningful threshold for relevance, a hybrid inflection point detection method is applied. Two complementary methods are combined:
  • Similarity-jump approach: Detects the point where cosine similarity values exhibit a sharp decline, marking the transition between relevant and irrelevant tools.
  • Kneedle algorithm: Detects inflection points in the similarity distribution by analyzing the deviation between the normalized cumulative similarity curve and the diagonal. The point of maximum deviation is chosen as the cutoff threshold for selecting relevant tools.
Step 3: Dynamic cutoff and final section
After both methods, the final retained tool count $N$ is set to the smaller of the two candidate sizes:
$N=\min(N_{\text{jump}},N_{\text{kneedle}}),$ (1)
where $N<M$. If $N<10$, the system supplements the set by selecting the top-10 most similar tools.
Empirical observations in real-world scenarios showed that overly small $N$ values restricted candidate diversity, whereas excessively large $N$ introduced semantic noise and redundant retrievals. Thus, setting $N=10$ achieved an optimal balance between precision and efficiency.
The refined tool set is passed to the LLM, where vector-based retrieval and inflection-point filtering together ensure accurate, efficient access while minimizing irrelevant calls and maintaining broad task coverage.
3.3 Hierarchical Memory Management
To address the challenge of context expansion in multi-turn dialogues, this module adopts a layered memory mechanism, which integrates fine-grained dialogue-level message alignment and coarse-grained session-level summarization.
(1) Dialogue-level memory alignment: Each round of interaction turn comprises of a sequence of message types: HumanMessages, AIMessages, and optionally ToolMessages, depending on whether external tools are invoked during the turn. The structure of the dialogue history is preserved as follows:
  • Without tool calls: The memory typically consists of two messages: one HumanMessage and one corresponding AIMessage.
  • With tool calls: the message flow becomes more complex. If $k$ tools are sequentially invoked in a single turn, the memory sequence contains $2+2k$ entries: one HumanMessage, $(k+1)$ AIMessages and $k$ ToolMessages. Concretely, $k=1\rightarrow 4$ messages (Human, AI, Tool, AI); $k=2\rightarrow 6$ messages (Human, AI, Tool, AI, Tool, AI); $k=3\rightarrow 8$ messages, and so on.
To maintain structural consistency and semantic coherence across turns, this module performs pairwise alignment of adjacent messages in the diaglogue traces. This alignment process detects and rectifies disruptions in the expected message sequence, particularly those arising from partial or failed tool executions (e.g., user-initiated cancellations, API failures, or LLM output errors) that result in missing or empty ToolMessages. In such cases, the system inserts semantically appropriate placeholder entries or reconstructs the absent information through contextual inference. Furthermore, to mitigate data loss caused by system faults or token-length limitations during tool invocation, the mechanism employs a backfilling strategy to recover truncated or omitted messages. By reconstructing these missing components based on available context and execution logs, the module ensures the integrity of the canonical interaction pattern: Human → AI (action proposal) → Tool → AI (response synthesis).
Collectively, these operations preserve the temporal and functional continuity of the dialogue state, thereby enhancing robustness and reliability in extended, tool-augmented conversations.
Figure 4: Message Summarization Process.
(2) Session-level memory optimization: When the cumulative number of messages within a session exceeds a predefined threshold $K$, the system triggers a context summarization mechanism to compress the historical dialogue state, thereby mitigating excessive token consumption and maintaining inference efficiency. Rather than employing naive truncation, which risks loss of critical semantic information, our framework implements a structured summarization strategy designed to preserve conversational coherence and intent continuity.
Specifically, the system identifies the segment for summarization as the subsequence extending from the earliest HumanMessage $H_{0}$ up to (but excluding) the second-to-last HumanMessage in the current history. This ensures that the most recent interactive context—containing the latest user input and subsequent system responses—remains unsummarized and fully accessible to support ongoing reasoning and tool invocation. For instance, given a message sequence (as shown in Figure 4), when the total message count exceeds capacity $K$, the segment from $H_{0}$ to $A_{5}$ is selected for summarization. The resulting summary is stored in the session state under state[”summary”] and subsequently integrated into the dialogue context as a SystemMessage. This summarized entry replaces the original message block in the history, effectively reducing memory footprint while retaining high-level semantic content.
During model interaction, the summary message is prepended to the active dialogue history, ensuring that salient contextual information is preserved at the beginning of the input sequence. This placement optimizes its retention under attention mechanisms and supports coherent continuation of long-horizon conversations.
By integrating fine-grained dialogue-level alignment with coarse-grained session-level summarization, the proposed memory architecture achieves robust, scalable context management. It effectively balances the trade-off between contextual fidelity and computational efficiency, minimizing redundancy while preserving task-critical intent across extended multi-turn interactions.
4 Experimental Design and Result Analysis
We present a evaluation of the proposed optimization modules of adaptive prompt generation, tool orchestration, and memory management, then integrated into several variants of ReAct-style autonomous agent frameworks. The primary objective is to evaluate how these components enhance procedural fidelity and semantic quality in complex, tool-augmented, multi-turn interactions, including token consumption, enabling a comprehensive evaluation of performance and resource utilization in real-world scenarios.
4.1 Evaluation Metrics and Assessment Criteria
The evaluation system consists of the following three parts, as shown in Figure 5:
Figure 5: Comprehensive agent evaluation framework for Procedural, Semantic, and Efficiency Dimensions.
4.1.1 Execution fidelity
Examines the agent’s behavioral precision and procedural consistency during task execution, focusing on the sequence of tool invocations and the degree of task completion.
This assessment goes beyond static output quality by examining the rationality and correctness of the agent’s operational trajectory, focus on how it is achieved. As such, it better aligns with the functional requirements of autonomous agents in real-world, dynamic environments where correct behavior depends on both accurate decisions and proper execution order.
To quantify execution fidelity, we introduce a structured evaluation framework based on four complementary metrics, collectively termed the 4T metrics, to assess critical aspects of tool-mediated reasoning and action:
  • Task Completion Rate (TCR): Proportion of tasks where all required tools are invoked in correct sequence:
    $\text{TCR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[R_{i}\subseteq L(i)\right],$ (2)
    where $N$ is the total number of test tasks, $L(i)$ is the actual tool sequence invoked by the agent on task $i$, and $R_{i}$ is the human-annotated reference sequence. The indicator function $\mathbb{I}[\cdot]$ returns 1 if $R_{i}$ is a subsequence of $L(i)$, and 0 otherwise.
  • Task Failure Rate (TFR): Percentage of tasks with no meaningful execution due to empty tool calls or runtime/system errors:
    $\text{TFR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[L(i)=\emptyset\lor\text{Error}(i)\right].$ (3)
    • No tools were invoked: $L(i)=\emptyset$
    • Runtime errors (e.g., recursion limit, tool crash)
    • Agent or system exceptions
  • Task Incompletion Rate (TIR): TIR measures the rate of partially completed tasks due to mistakes in tool selection or sequence, excluding runtime errors. These include missing required tools, misordered invocation, or calling unrelated tools.
    $\text{TIR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[R_{i}\not\subseteq L(i)\land L(i)\neq\emptyset\right].$ (4)
  • Task Performance Score (TPS): Fine-grained measure considering correct, wrong, and missing tools:
    $\begin{gathered}\text{TPS}(i)=\frac{C_{i}}{C_{i}+\lambda_{w}W_{i}+\lambda_{m}M_{i}},\\ \text{TPS}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}\text{TPS}(i),\end{gathered}$ (5)
    where $C_{i}$, $W_{i}$, and $M_{i}$ denote the number of correctly used, wrongly invoked, and missed tools, respectively. The penalty weights $\lambda_{w}$ and $\lambda_{m}$ default to 1 unless otherwise specified.
Together, TCR, TFR, TIR, and TPS provide a comprehensive assessment of procedural fidelity, capturing both task completion and the rationality of tool usage.
4.1.2 Output quality
This metric evaluates the semantic, structural, and contextual consistency between the agent’s response and the reference answer.
To achieve scalable, fine-grained evaluation beyond lexical overlap, we adopt an LLM-based assessment framework using models such as Qwen-3 and DeepSeek as intelligent evaluators. This approach overcomes the limits of traditional metrics (e.g., BLEU, ROUGE) by jointly considering the response, reference, and contextual metadata (user query, prompt, and tool descriptions).
The evaluator scores five dimensions, denoted as CRCFF metrics:
  • Correctness: Whether the response accurately reflects factual content without errors or distortions.
  • Relevance: Whether it directly addresses user intent and aligns with task context.
  • Completeness: Whether it covers all required information without omissions.
  • Fluency: Whether the text is coherent, natural, and grammatically correct.
  • Faithfulness: Whether it remains logically consistent and free from hallucinated details.
Each is rated from 0–10, enabling interpretable, multidimensional evaluation of response quality.
Table 1: Comparison of APIGen and Jenius-bench.
Aspect APIGen Jenius-bench
Dialogue Turn Single Multi
Data Source LLM-generated Human + Automatic
Verification Format / Execution / Semantics Trace + Review
Task Type Simple calls Chained reasoning
Tool Info Basic meta Contextual MCP
Memory None Cross-turn
Realism Abstracted Real-world
Scale 60K / 21 cats. 850 / 38 cats.
Focus Call accuracy Reasoning / Planning
4.1.3 Token consumption
Measures the computational input and output overhead of the agent by quantifying the number of input tokens fed to the LLM and generate output tokens.
This metric captures the cumulative context length, including user queries, system prompts, tool descriptions, and historical dialogue states, and serves as an indicator of inference cost, memory usage, and scalability.
4.2 Benchmark Task Design and Dataset Construction
To comprehensively evaluate the capabilities of autonomous agents in diverse operational contexts, we conducted experiments on two complementary benchmark datasets: one widely adopted public benchmark and one newly curated, real-world task-oriented dataset. These datasets differ significantly in their dialogue structure, tool complexity, metadata richness, and annotation rigor, enabling a multi-faceted assessment of agent performance across controlled and realistic scenarios.
4.2.1 APIGen: A Single-Turn Tool-Use Benchmark
APIGen Liu et al. (2024) evaluates models’ ability to interpret natural language queries into structured API calls. It contains 60K annotated samples across 21 categories (3.6K APIs, 2.3 parameters per call). Each instance includes a user query, candidate APIs, and a ground-truth function call.
In our evaluation, we sample 800 instances from APIGen to ensure computational tractability while preserving task diversity. However, a key limitation of APIGen is its restriction to single-turn interactions: each instance represents an isolated tool-calling decision without contextual dependencies or multi-step reasoning. Furthermore, the API descriptions are minimal and syntactic in nature, lacking rich semantic annotations or behavioral specifications. To amplify the challenge and better stress to test our retrieval-augmented architecture, we extend the original candidate tool list for each query by randomly injecting 100 irrelevant tools sampled from its API pool, increasing the average candidate count. This augmented setup intensifies the need for accurate semantic matching and efficient top-$M$ filtering, thereby highlighting the effectiveness of our dynamic tool selection mechanism under high-noise conditions.
4.2.2 Jenius-bench: A Multi-Turn, Semantically Enriched Task Dialogue Dataset
To address the limitations of synthetic and single-turn benchmarks, we introduce Jenius-bench, a novel, task-oriented dialogue dataset specifically constructed to evaluate advanced agent systems in realistic, interactive environments. A detailed comparison between APIGen and Jenius-bench is provided in Table 1. Comprising 850 high-quality, human-curated samples, Jenius-bench covers a broad range of application domains—including travel planning, air and train ticket booking, web search, weather inquiry, website generation, and academic paper retrieval—spanning 38 distinct tool categories.
Crucially, Jenius-bench supports multi-turn dialogues with complex state evolution, where each interaction step includes detailed execution traces: the invoked tool (implemented as a fully encapsulated MCP-compliant module), its input arguments, the actual output returned, and the agent’s subsequent response. Unlike APIGen’s shallow metadata, tools in Jenius-bench are associated with rich semantic descriptions, including purpose, preconditions, expected outcomes, and usage examples, enabling deeper intent understanding and more robust orchestration.
Moreover, all trajectories in Jenius-bench are derived from real user-agent interactions and undergo rigorous manual review and annotation by domain experts to ensure correctness, coherence, and faithfulness. The final responses and execution paths serve as gold-standard references, making the dataset particularly suitable for evaluating not only output quality but also procedural fidelity and error recovery in long-horizon tasks.
4.3 Comparative Agent Configuration and Ablation Design
To systematically evaluate the individual and cumulative contributions of the proposed optimization modules, we conducted controlled ablation studies across four progressively enhanced agent configurations. Each variant shares a consistent architectural backbone and inference protocol, differing only in the inclusion of specific components, thereby enabling fine-grained attribution of performance changes to targeted improvements.
All agents are evaluated on both APIGen and Jenius-bench using our assessment framework: process fidelity is measured using the 4T metrics, while outcome quality is assessed via the CRCFF metrics. Additionally, total token consumption is reported to evaluate computational efficiency.
The evaluated agents are defined as follows:
  • Base: A standard ReAct-style agent implementing the observe–think–act loop, generating reasoning traces and tool calls dynamically. It serves as the canonical baseline for modular agent systems.
  • B-P: Incorporates adaptive prompt generation, dynamically reformulating system prompts based on task intent classification to improve contextual alignment.
  • B-PT: Further extends B-P with context-aware tool orchestration, retrieving relevant tools via dense semantic matching using Qwen3-Embedding.
  • Jenius: The complete system combining all three modules, adding hierarchical memory management with dialogue- and session-level summarization for stable long-horizon reasoning.
Table 2: Execution Fidelity Results on APIGen
Model TCR TFR TIR TPS
Base 0.8150 0.1800 0.0050 0.8150
B-P 0.8275 0.1675 0.0050 0.8275
B-PT 0.8375 0.1587 0.0038 0.8375
Jenius 0.8500 0.1362 0.0138 0.8500
Table 3: Execution Fidelity Results on Jenius-bench
Model TCR TFR TIR TPS
Base 0.5659 0.0329 0.4012 0.5968
B-P 0.7271 ($\uparrow$16%) 0.0859 0.1871 0.7491
B-PT 0.7494 0.0718 0.1788 0.7740
Jenius 0.7647 0.0753 0.1600 0.7847
4.4 Experimental Results and Comparative Analysis
4.4.1 Execution fidelity - 4T testing performances
Table 4: Output Qualit Results on Jenius-bench Evaluated by Qwen and DeepSeek
Evaluator Model Correctness Relevance Completeness Fluency Faithfulness
Qwen Base 0.6741 0.8951 0.7722 0.9294 0.7919
B-P 0.7259 ($\uparrow$5.2%) 0.9204 0.7873 0.9563 0.8390
B-PT 0.7572 0.9380 0.8027 0.9763 0.8560
Jenius 0.7580 0.9447 0.8088 0.9771 0.8766
DeepSeek Base 0.7890 0.9245 0.7898 0.9546 0.8291
B-P 0.8310 ($\uparrow$4.2%) 0.9357 0.8153 0.9613 0.8654
B-PT 0.8330 0.9428 0.8174 0.9711 0.8652
Jenius 0.8350 0.9400 0.8143 0.9686 0.8636
The results demonstrate distinct patterns between the two datasets. On the public APIGen benchmark, all models perform similarly, with the Base model already reaching a high task completion rate (TCR = 0.8150). Subsequent optimizations (B-P, B-PT, and Jenius) yield only marginal improvements up to 0.8500, reflecting that APIGen mainly assesses pattern matching within familiar tool templates rather than genuine reasoning or generalization.
In contrast, on the Jenius-bench dataset (Table 3), the benefits of modular optimization are substantial. B-P improves TCR from 0.5659 to 0.7271 and reduces redundant invocations (TIR) by 21%. B-PT further enhances accuracy (TCR = 0.7494, TPS = 0.7740), while the fully integrated Jenius model achieves the best overall performance (TCR = 0.7647, TPS = 0.7847), exhibiting coherent multi-turn tool orchestration. Overall, the saturation observed on APIGen can be attributed to extensive pre-training coverage, whereas Jenius-bench underscores the robustness and adaptability of our modular agent optimization framework.
Figure 6: Token consumption of evaluated agents on APIGen and Jenius-bench.
4.4.2 Output quality - CRCFF evaluation metrics
We evaluate generated responses using two independent LLM-based evaluators (Qwen and DeepSeek) across five textual quality dimensions: correctness, relevance, completeness, fluency, and faithfulness (CRCFF). As shown in Table 4, the evaluation is conducted solely on Jenius-bench, since APIGen employs virtual API stubs that lack real textual outputs for reliable quality analysis. Jenius-bench, in contrast, features realistic multi-turn dialogues where reasoning traces and tool results jointly determine response quality. The results from both evaluators reveal consistent trends: the Adaptive Prompt Generation module significantly enhances relevance and fluency by producing more stable, context-aware prompts; the Context-Aware Tool Orchestration module improves correctness and completeness through precise semantic alignment and richer tool utilization; and Hierarchical Memory Management strengthens faithfulness by preserving contextual coherence in long-horizon tasks. Overall, the fully integrated Jenius Agent achieves the best CRCFF performance, with an 8–10% improvement in correctness and fluency over the baseline, demonstrating the complementary effectiveness of the three optimization modules in producing coherent, accurate, and faithful responses.
4.4.3 Token consumption analysis
We compare token efficiency of all agents on APIGen and Jenius-bench (Figure 6), focusing on tokens used during tool invocations. In APIGen, only tool-call tokens are measured, as its virtual API stubs produce no real outputs.
Across both datasets, results show a clear downward trend. On APIGen, total tokens drop from 9.96M in Base to 2.41M in B-PT, with the full Jenius model maintaining similar efficiency (2.46M). On Jenius-bench, usage decreases from 9.27M to 3.65M, a reduction of over 60% compared with the baseline. These results highlight the cumulative efficiency gains brought by our modular design.
Specifically, Adaptive Prompt Generation reduces redundant reasoning, Context-Aware Tool Orchestration minimizes unnecessary tool calls, and Hierarchical Memory Management preserves coherence without increasing overhead. Overall, the consistent decline in token usage from Base to Jenius confirms that each module contributes complementary benefits, enabling efficient and interpretable agent behavior for real-world deployment.
5 Real-World Deployment and Experience
This section presents the real-world deployment and operational insights of our anonymous agent system of Jenius in a production environment. We outline the deployment context, summarize key usage metrics over a four-month period, and distill practical lessons learned from operating the system at scale.
5.1 Deployment Context
The Jenius system was initially rolled out through a gray-scale release on May 21, 2025, followed by full public deployment on May 25, 2025. Since then, the system has undergone continuous optimization guided by the methodology introduced in prior sections.
Hosted on Alibaba Cloud, the architecture adopts a multi-node distributed design to ensure modularity, scalability, and fault isolation. Key components include: the Jenius Agent Engine for core reasoning and tool orchestration, several dedicated MCP Tool Servers for secure function execution, an external LLM Service for language generation, and Langfuse as the observability and logging infrastructure for tracing interactions and debugging.
All services are containerized and orchestrated using Kubernetes with integrated CI/CD pipelines, enabling automated deployment and version management.Inter-service communication is secured via mutual TLS, ensuring data integrity and confidentiality. This modular design enables independent scaling and monitoring of each subsystem while maintaining end-to-end privacy and reliability.
5.2 Scale and Usage Patterns
  • During the first month of public operation, the system achieved a daily active user (DAU) count of 42 and a monthly active user (MAU) count of 1,277, indicating early but concentrated adoption.
  • User engagement was characterized by high interaction frequency: on average, each active user initiated 2.06 conversation sessions per day, with each session consisting of 2.4 turns (i.e., message exchanges between user and agent). Furthermore, every turn involved an average of 1.15 tool calls, reflecting the system’s reliance on external functions to fulfill complex requests.
  • Geographically, usage spanned 34 countries, with the majority of traffic originating from Mainland China (97.07%), followed by the United States (1.26%) and Japan and Malaysia (combined <1<1%). Within China, users were distributed across nearly 300 cities, with the highest concentrations in Beijing (8.67%), Suzhou (8.03%), Shanghai (6.01%), and Guangzhou (5.39%). This distribution suggests strong initial traction in major technology and business hubs.
Despite the relatively small user base, the depth of interaction demonstrates meaningful utility and sustained user engagement.
5.3 Lessons Learned
Real-world deployment of the Jenius Agent provides critical insights into the design and optimization of large language model LLM-based agents.
  1. 1.
    Balancing Internal Knowledge and External Tool Use: We observed that relying on tool calls for every query, especially those answerable from the model’s internal knowledge, leads to unnecessary computational overhead and increased latency. By strategically leveraging the LLM’s pretrained knowledge for common or factual queries (e.g., definitions, general knowledge), we significantly reduced token consumption and API costs while improving response speed. This hybrid approach enhances efficiency without compromising accuracy.
  2. 2.
    Prudent Tool Invocation with Semantic Fidelity: A recurring issue was the generation of spurious tool calls: either invoking non-existent tools or fabricating parameters not grounded in user input. To mitigate this, we implemented stricter prompting strategies and post-generation validation checks to ensure that tool invocations are both necessary and semantically justified. This principle of “think before calling” is essential for maintaining system reliability and trustworthiness.
  3. 3.
    Explainability in Multi-Step Reasoning: For complex tasks requiring several sequential tool calls, users often found the interaction confusing due to the lack of intermediate explanations. We found that providing concise, human-readable reasoning traces, such as summarizing the plan or justifying each step, greatly improves transparency and user comprehension. Explicit planning and explanation are therefore crucial for usability in long-horizon tasks.
  4. 4.
    Error Handling and Retry Behavior: The default retry mechanism for failed tool calls led to excessive resource consumption and degraded user experience, particularly when errors were irrecoverable (e.g., network timeouts, invalid inputs). Blind retries not only wasted compute resources but also caused frustrating delays. We addressed this by introducing adaptive retry policies based on error types and user feedback signals, reducing redundant executions by over 60% in subsequent iterations.
  5. 5.
    Specialized Handling for High-Cost or Fragile Tools: Certain tools require domain-specific safeguards. For instance, the ’url_reader’ tool initially failed to extract URLs due to mismatched patterns in user input; this was resolved by incorporating robust regular expression matching. Similarly, long-running operations such as PPT or web application generation were prone to being triggered inappropriately (e.g., interpreting “generate a fitness plan” as a request for a PowerPoint). We introduced intent disambiguation layers and explicit confirmation prompts to prevent unintended activation of expensive or irreversible actions.
These findings demonstrate that our closed–loop optimization framework is not only academically sound but also practically impactful, providing a blueprint for integrating LLM–based agents into large–scale industrial systems.
6 Discussion and Conclusion
This work enhances autonomous agent performance through three complementary modules—adaptive prompt generation, context-aware tool orchestration, and hierarchical memory management. These modules jointly improve reasoning stability, tool invocation accuracy, and context retention, leading to consistent gains across text quality, execution fidelity, and token efficiency. The evaluation framework combines procedural, semantic, and efficiency metrics, revealing strong overall improvements while also highlighting limitations such as incomplete capture of hidden reasoning steps and the existence of multiple valid tool-use paths. Despite these challenges, large-scale deployment confirms the reliability and generalizability of the proposed framework.
Future work will focus on more flexible and outcome-oriented evaluation, accommodating diverse reasoning trajectories and tool sequences without penalizing valid outputs. Integrating user satisfaction, decision-making cost, and latency into evaluation will further align metrics with real-world performance. Operationally, adaptive strategies that reconfigure modules based on dynamic task contexts and multi-agent collaboration for distributed problem-solving represent key directions. The resulting system, Jenius, demonstrates stable performance improvements, efficient token utilization, and high user satisfaction in real-world use, providing both a validated empirical foundation and a practical blueprint for scalable, production-ready autonomous agent systems.